# Securing Federated Learning Against Novel and Classic Backdoor Threats During Foundation Model Integration

### v1

@@ Abstract
Federated learning (FL) enables decentralized model training while preserving privacy. Recently, integrating Foundation Models (FMs) into FL has boosted performance but also introduced a novel backdoor attack mechanism. Attackers can exploit the FM’s capabilities to embed backdoors into synthetic data generated by FMs used for model fusion, subsequently infecting all client models through knowledge sharing without involvement in the long-lasting FL process. These novel attacks render existing FL backdoor defenses ineffective, as they primarily detect anomalies among client updates, which may appear uniformly malicious under this attack. Our work proposes a novel data-free defense strategy by constraining abnormal activations in the hidden feature space during model aggregation on the server. The activation constraints, optimized using synthetic data alongside FL training, mitigate the attack while barely affecting model performance, as the parameters remain untouched. Extensive experiments demonstrate its effectiveness against both novel and classic backdoor attacks, outperforming existing defenses while maintaining model performance.
@@
Index Terms—  Backdoor Defense, Federated Learning, Foundation Models, Adversarial Machine Learning
1 Introduction
@@
Federated learning (FL) enables collaborative model training across decentralized institutes or devices, enhancing data privacy and security with applications in healthcare [1, 2], finance [3], and IoT [4]. Recently, integrating Foundation Models (FMs), such as GPT series, LLaMA series, and Stable Diffusion, introduces new dynamics to FL. Pre-trained on large datasets, FMs excel in tasks from natural language processing to vision data generation and recognition, augmenting FL by improving performance through knowledge distillation [5, 6] or enabling client knowledge sharing with synthetic data generation [7, 8].
However, incorporating FMs into FL systems also introduces new complexities and exacerbates existing threats. The inherent vulnerabilities of FMs create additional attack vectors through their interaction with FL. Recent studies [9, 10] show that FMs’ in-context learning (ICL) capabilities enable attackers to exploit novel attacks against FL. Specifically, an attacker can operate externally to the FL process without maintaining its presence throughout its duration. For instance, in a typical FM-integrated FL framework, a large language model (LLM) generates synthetic data for aggregating client model parameters. An attacker can embed a backdoor into the LLM-generated synthetic dataset using malicious prompts during inference. This backdoor is subsequently transmitted to all FL clients during knowledge sharing, thereby undermining their integrity.
@@
Existing FL backdoor defense methods, which typically detect anomalies among clients during client model aggregation, are ineffective against the novel attacking mechanism [11, 12, 13, 14, 15]. Since all client updates appear uniformly malicious due to compromised synthetic data, no detectable anomalies arise. Moreover, many general machine learning (ML) backdoor defenses are unsuitable for FL. Some require access to clean, real datasets [16, 17], conflicting with FL’s data isolation principle, while others are time-intensive [18], making them impractical for continuous FL training cycles.
Given the novel threats from ML advances, it is critical to develop an effective defense strategy that addresses these challenges within the FL framework. Since malicious updates could potentially originate from the majority of clients rather than being limited to outliers, a feasible method is to periodically constrain potential malicious updates from all clients on the server. Studies [19, 20, 17] show that backdoor attacks often cause abnormally large activations in the hidden layers of the compromised model. To mitigate this, one could impose upper bounds on the internal activations following the aggregation of client updates. These constraints aim to keep activations within reasonable limits, effectively mitigating the risk of novel backdoor attacks. The optimal upper bounds can be adjusted to balance model performance on certain datasets, e.g., the synthetic dataset. Moreover, the proposed approach suits FL better than general ML backdoor defenses. It uses synthetic data for constraint optimization, preserving local data independence, and periodic adjustment of bounds is more efficient than parameter tuning or trigger estimation strategies.
@@
In summary, this paper presents the following contributions: (1) We introduce the first data-free defense strategy against the novel backdoor attacks arising from the integration of FMs into FL. (2) Extensive experiments in diverse FL scenarios validate the effectiveness of our defense against both novel and classic backdoor attacks, which are unified within the proposed defense framework.
2 Related Work
Vulnerabilities Introduced by FM integrated FL: The interaction between FMs and FL enhances both domains [21] but also introduces new attack vectors [22]. A common use of FM-integrated FL is generating synthetic data for model pre-training [23], knowledge transfer [6], and model aggregation [7]. However, FMs, especially LLMs with ICL ability, are vulnerable to inference-time poisoning [24, 25, 26]. Recent studies [9, 10] reveal a novel attack mechanism exploiting this capability. The attacker uses malicious prompts to embed a backdoor in LLM-generated synthetic data, which is then transmitted to all FL clients during model aggregation, compromising their integrity.
@@ 2 Related Work
FL Defenses: The existing FL defenses[11, 27, 12, 13, 14, 15] primarily target client-originated threats and offer limited protection against this attack approach. [11] applies a fixed norm threshold to client updates in FL. [27] combines norm thresholding of client updates with the injection of Gaussian noise into the aggregated global model at the server end. [12, 13, 14, 15] selects the most reliable gradient update from all participants to counter adversarial or faulty updates. [28] employs federated aggregation of neuron activation values to prune the least active neurons.
3 Methodology
3.1 FM integration in FL
Our work follows existing FM-integrated FL frameworks, such as [21, 23]. An LLM on the server generates synthetic data for model aggregation, including text data and prompts guiding other FMs to produce data in formats like images. Combined with classic frameworks like [7], the FM-integrated FL cycle involves local client training, update uploads, “Ensemble Distillation” based model fusion using the synthetic data, and aggregated parameters distribution.
3.2 Threat Model and Assumptions on Defender
@@
Attacker’s knowledge: The attacker lacks access to the local training set and process, distinguishing it from traditional backdoor attacks. Instead, they exploit access to the server’s LLM queries to insert malicious instructions, specifying the trigger, target class, and demonstrations that show how the attack is activated.
Attacker’s goals: The attacker aims to (1) direct the LLM and other FMs to generate synthetic datasets with a percentage of backdoor poisoned samples, and (2) leveraging (1), propagate the backdoor to all client models in FL, causing the final model to misclassify triggered inputs to the target class while maintaining high performance on clean samples.
Attack scenario: Given the diversity of LLMs and the goal of improving outputs, users may rely on third-party services for API integration and prompt engineering. In this case, the attacker could be the service provider, thereby accessing the user’s queries.
@@
Defender’s goals: The defender aims to counteract backdoor attacks during FL training, ensuring the final model matches clean model performance. Specifically, it should classify backdoor-embedded inputs correctly while maintaining high accuracy on clean samples.
3.3 Method
@@
Existing defenses face challenges with the novel attack: (1) Most FL defenses [11, 12, 14, 15, 13] detect anomalies in client updates but are ineffective against novel attacks where malicious updates may come from the majority of clients. (2) General ML defenses need clean, real datasets, which contradicts the decentralized nature of FL [16, 17]; (3) The ongoing nature of FL makes repeated use of time-consuming ML defenses like trigger estimation [18] impractical.
@@
Given the unique attack mechanism and limitations of current defenses, a feasible strategy is to periodically constrain all client (malicious) updates on the server. Studies[19, 20, 17] show backdoor attacks often cause abnormally large activations in hidden layers. Hence, malicious updates can be mitigated by setting upper bounds on internal activations after client update aggregation, with the bounds optimized to maintain model performance on certain datasets. To address challenge 2, upper bounds optimization can be performed using the (possibly poisoned) synthetic dataset, preserving the independence of local data and training in FL. This method minimizes impact on the model’s classification ability without altering underlying parameters. Additionally, periodically adjusting these bounds is more efficient and practical than the extensive parameter tuning or trigger estimation used in other contexts, addressing challenge 3.
@@
Now we elaborate on the server-side defense in FL. It is implemented after the “Ensemble Distillation” based model fusion using the synthetic data $D_{\text{syn}}$ [^1] The processes of client local training and “Ensemble Distillation” based model fusion, which remain unaltered, are not covered here. For details, refer to [7]. Let $g(\cdot|\theta)$ denote the $L$-layer model with parameters $\theta$, derived from the model fusion at FL round $t$ [^2] For clarity, we omit $t$ in our notation.. We define the set $\mathcal{K}\subseteq\{1,\cdots,L\}$ to include all layers where bounds are applied, and the set $\mathcal{Z}=\{\mathbf{z}_{k}\in\mathbb{R}^{n_{k}}|k\in\mathcal{K}\}$ to encompass all vectors used to constrain activations at layer $k$ within $\mathcal{K}$. Then the logit function for any class $c\in\mathcal{C}$ and any input $\mathbf{x}\in\mathbb{R}^{n_{0}}$ with activation bounds $\mathcal{Z}$ can be written as
$\overline{g}_{c}(\mathbf{x;\mathcal{Z}})=\mathbf{w}_{c}^{T}(\tau_{L}\circ\sigma_{L}\circ\cdots\circ\tau_{1}\circ\sigma_{1}(\mathbf{x})),\vspace{-0.08in}$
@@
where $\sigma_{l}:\mathbb{R}^{n_{l-1}}\to\mathbb{R}^{n_{l}}$ is the composition of the weight and activation function at the $l$-th layer of the model, $\mathbf{w}_{c}^{T}\in\mathbb{R}^{n_{L}}$ is the weight vector associated with class $c$, and $\tau_{l}(\cdot)$ applies upper bounds. For layer $l\in\mathcal{K}$, $\tau_{l}(\mathbf{v})=\min\{\mathbf{v},\mathbf{z}_{l}\}$, applied element-wise to the vectors; otherwise, $\tau_{l}(\mathbf{v})=\mathbf{v}$.
To ensure effective defense, activation upper bounds are minimized to allow only legitimate activations through while filtering out malicious ones. Since the defender lacks knowledge of attack triggers and target classes, the optimization relies solely on clean samples. Hence, one can minimize the upper bounds so that the classification accuracy on clean samples exceeds a certain threshold.
However, in the defense scenario considered here, the defender lacks access to clean, real data. Consequently, we adjust the optimization problem to align the logits of synthetic data - with bounds applied - with those of the original model. This approach helps preserve the effectiveness of the defense and the model’s functionality as much as possible by: (1) optimizing the bounds based on the model’s relative performance on the synthetic data, thus avoiding reliance on absolute (possibly poor) classification accuracy on the synthetic data; and (2) keeping the model parameters unchanged. Motivated by the above, we optimize the upper bounds by minimizing the following Lagrangian function:
$H(\mathcal{Z},\lambda)=\sum_{\mathbf{x}\in D_{\text{syn}},c\in C}[\overline{g}_{c}(\mathbf{x;\mathcal{Z}})-g_{c}(\mathbf{x})]^{2}+\lambda{\sum_{l\in\mathcal{K}}{\lVert\mathbf{z}_{l}\rVert}_{2}}\vspace{-0.1in}$
@@
For an effective defense, we dynamically adjust the Lagrangian multiplier $\lambda$ during FL. Let the threshold $\Delta\pi$ represent the allowable drop in classification accuracy on synthetic data. A lower threshold preserves performance but may miss constraining misclassifications, while a higher threshold may overly constrain. If the accuracy drop is below $\Delta\pi$, we increase $\lambda$ by $\alpha$ to tighten bounds; if it exceeds $\Delta\pi$, we decrease $\lambda$ to relax the constraints.
Note that the classic backdoor attack is a special case of the novel attack in our defense framework. Activation bounds also suppress malicious updates from few compromised clients. Moreover, after model fusion and activation bounds optimization, only the model parameters are distributed to clients. Separating bounds from local training prevents compromised clients from adapting to the bounds and thus evading the defense. Additionally, only a few optimization iterations are needed to ensure effective defense, efficiently securing the FL training process.
4 Experiment
4.1 Experimental Setup

### v2

# Securing Federated Learning against Backdoor Threats with Foundation Model Integration

@@ Abstract
Federated Learning (FL) enables decentralized model training while preserving privacy. Recently, the integration of Foundation Models (FMs) into FL has enhanced performance but introduced a novel backdoor attack mechanism. Attackers can exploit FM vulnerabilities to embed backdoors into synthetic data generated by FMs. During global model fusion, these backdoors are transferred to the global model through compromised synthetic data, subsequently infecting all client models. Existing FL backdoor defenses are ineffective against this novel attack due to its fundamentally different mechanism compared to classic ones. In this work, we propose a novel data-free defense strategy that addresses both classic and novel backdoor attacks in FL. The shared attack pattern lies in the abnormal activations within the hidden feature space during model aggregation. Hence, we propose to constrain internal activations to remain within reasonable ranges, effectively mitigating attacks while preserving model functionality. The activation constraints are optimized using synthetic data alongside FL training. Extensive experiments demonstrate its effectiveness against both novel and classic backdoor attacks, outperforming existing defenses.
Index Terms:
@@
Backdoor Defense, Federated Learning, Foundation Models, Adversarial Machine Learning
I Introduction
@@
Federated learning (FL) enables models to be collaboratively trained across multiple decentralized institutes or devices. This method significantly enhances data privacy and security and has a wide range of applications, including healthcare[22, 32], finance[20], and IoT[21]. Recently, the integration of Foundation Models (FMs) like the GPT series [5], LLaMA [28], and Stable Diffusion [24] has introduced new dynamics to FL[39]. These models, pre-trained on vast and varied datasets, show remarkable proficiency in various tasks from natural language processing to vision data generation and recognition. By leveraging their extensive pre-trained knowledge, FMs can augment FL systems either by enhancing performance through knowledge distillation [13, 37] or by facilitating client knowledge sharing with synthetic data generation [18, 1].
However, incorporating FMs into FL systems (FM-FL) introduces new complexities and exacerbates existing threats. Recent studies [15, 16] have identified a novel backdoor mechanism in FL arising from the interaction between FL and FM. Consider a classic FM-FL scenario where the FM on the server generates synthetic data that mirrors the client data distribution for use in global model aggregation. An attacker can exploit FM vulnerabilities to embed a backdoor into the synthetic data. This backdoor is then planted into the global model during the aggregation stage using the poisoned synthetic dataset. Subsequently, the backdoor is transmitted to all FL clients as the poisoned global model is distributed, thereby compromising the integrity of the entire system.
@@
Existing FL backdoor defense methods cannot address the novel attack due to its fundamentally different attack mechanism compared to the classic one. In the classic attack, a few clients are compromised first, and the global model is subsequently infected through client model fusion. To mitigate such attacks, existing methods detect anomalies in client updates and exclude them from model aggregation [26, 4, 14, 35, 23]. In contrast, the novel attack introduced by the integration of FMs compromises the global model first through the use of poisoned synthetic data, which then propagates the backdoor to all client models. Since the local data remains untouched, all client updates appear similar, and no detectable anomalies arise. As a result, existing FL defenses are ineffective against this type of attack. Moreover, many general machine learning (ML) backdoor defenses are unsuitable. Some require access to clean, real datasets [17, 19], conflicting with FL’s data isolation principle, while others are time-intensive [29], making them impractical for continuous FL training cycles.
It is critical to develop an effective defense strategy that addresses both novel and classic backdoor threats within the FL framework. A common attack pattern in both types of attacks is the abnormally large activations in the hidden layers of the compromised model when processing manipulated inputs [30, 19]. To mitigate this, a feasible approach is to impose upper bounds on internal activations. These activation bounds aim to constrain activations to reasonable limits, effectively mitigating malicious behavior on poisoned samples while preserving the model’s functionality on normal samples. The optimization of these bounds leverages synthetic data, maintaining local data independence in FL. This process is performed after global model fusion on the server, with the original model parameters fixed, ensuring both functionality and defense efficiency. Hence, this method provides an effective defense against backdoor attacks while remaining well-suited to the FL framework.
@@
In summary, this paper presents the following contributions: (1) We propose the first data-free defense strategy that addresses both novel and classic backdoor attacks in FL with FM integration. (2) Extensive experiments across diverse FL scenarios validate the effectiveness of our unified defense framework against both types of backdoor attacks.
II Related Work
II-A Backdoor Attacks in FL
A backdoor attacker in FL aims to embed malicious behavior into the global model distributed to all clients. This backdoor behavior (e.g., misclassification to a specific target class) is triggered only by specific patterns embedded in input samples, while the model functions normally on clean inputs.
Classic Backdoor Threats: Classic backdoor threats primarily target the client side through techniques like data poisoning [27], local model poisoning [8], and attacks such as semantic and distributed backdoors [2, 31, 34]. For instance, attackers may inject poisoned samples into the local training datasets of compromised clients. These compromised local models then propagate the malicious model updates to the global model during server-side aggregation. With sufficient compromised clients and communication rounds, the global model is embedded with backdoor threats.
Novel Backdoor Threats: Recent studies [15, 16] have revealed a novel backdoor attack mechanism in FL that exploits the interaction between FMs and FL. Attackers leverage vulnerabilities in FMs [7, 25, 36] to embed backdoors into FM-generated synthetic data. This poisoned data is subsequently used for purposes such as model pre-training [38], knowledge transfer [37], or model aggregation [18]. The backdoor threat is then incorporated into the global model through the use of synthetic data and subsequently transmitted to all FL clients.
Classic vs. Novel Attack Mechanisms: Classic backdoor threats stem from a subset of compromised clients and propagate to the central server, with malicious updates identified as anomalies among client updates. In contrast, novel attacks originate on the server, embedding threats into the global model and affecting all clients uniformly, making anomalies undetectable.
II-B FL Backdoor Defenses
@@
Existing FL defenses [26, 9, 4, 14, 35, 23] primarily target client-originated threats and offer limited protection against novel attack approaches. These methods focus on identifying anomalies in client updates and excluding the updates from identified malicious clients during global model aggregation to achieve a robust global model. Specifically, [26] applies a fixed norm threshold to client updates in FL. [9] combines norm thresholding with the injection of Gaussian noise into the aggregated global model at the server. [4, 14, 35, 23] select the most reliable gradient updates from all participants to counter adversarial or faulty updates. [33] utilizes federated aggregation of neuron activation values to prune the least active neurons.
III Methodology
III-A Overview
FM integration in FL. Our work follows existing FM-integrated FL (FM-FL) frameworks, such as as those proposed in [39, 38]. The basic FM-FL cycle, as illustrated in Fig.1 and Algorithm1, consists of three key steps. Stage 1: Initialization. An FM is integrated into the server to generate synthetic data (e.g., text or image data) that mirrors the distribution of client-local data and is later used to fuse a global model. Stage 2: Client Update. Clients independently train their local models using private local data. Once trained, they upload their model parameters to the server for aggregation during the model fusion process. Stage 3: Server Global Model Fusion. The server aggregates the client model parameters using synthetic data as a carrier for client model information sharing. This process employs aggregation functions such as those proposed in [18, 13], which are applicable to various FL settings.
The proposed defense. To provide a unified defense strategy against both classic and novel backdoor attacks in FL, we propose Stage 4: Server Backdoor Mitigation, implemented on the server after the global model fusion. Regardless of whether the backdoor attacks are novel or classic, poisoned samples typically exhibit abnormal internal activations. Hence, the proposed defense optimizes an upper bound on the internal activations of the global model. This bound constrains the abnormal activations caused by poisoned samples while allowing the normal activations of clean samples to pass.
Stages 2–4 are repeated until FL converges. Throughout this process, we aim to minimize poisoning impact while maintaining effective training and aggregation.
III-B Threat Model and Assumptions on Defender
@@
Attacker’s knowledge: The attacker lacks access to the local training set and process, distinguishing it from traditional backdoor attacks. Instead, they exploit access to the server’s FM queries to insert backdoor instructions, specifying the trigger, target class, and demonstrations that show how the attack is activated [15, 16].
Attacker’s goals: The attacker aims to (1) guide the FMs to generate synthetic datasets containing backdoor-poisoned samples, and (2) leveraging (1), propagate the backdoor to all client models in FL, causing the final model to misclassify triggered inputs to the target class while maintaining high performance on clean samples.
@@
Defender’s goals: The defender aims to mitigate backdoor attacks during FL training, ensuring the final model matches clean model performance. Specifically, it should classify backdoor-embedded inputs correctly while maintaining high accuracy on clean samples.
III-C Server-side Backdoor Mitigation
Figure 1: Our defense strategy for the FM-FL framework.
@@ III-C Server-side Backdoor Mitigation
Existing defenses, initially proposed for classic backdoor attacks that arise from compromising several clients, face challenges with the novel backdoor attack. Challenge 1. Most FL defenses [26, 4, 35, 23, 14] detect anomalies in clients but are ineffective against novel attacks where malicious updates may come from the majority of clients. Challenge 2. General ML defenses [17, 19] need clean, real datasets. This contradicts the decentralized nature of FL, where no real data is available on the server. Challenge 3. The ongoing nature of FL makes repeated use of time-consuming ML defenses like trigger estimation [29] impractical.
@@
To address Challenge 1, we focus on the underlying attack pattern common to both novel and classic backdoor attacks in FL. Whether the attack originates from a few compromised clients or arises through the interaction between FM and FL, the poisoned global model exhibits the same malicious pattern: abnormally large activations in hidden layers triggered by poisoned inputs. This pattern has been observed in prior studies [30, 19]. Leveraging this shared malicious pattern, we propose transitioning from traditional client-level anomaly detection to activation-level anomaly detection, offering a unified perspective on backdoor defense. Malicious updates can be mitigated by imposing upper bounds on internal activations of the global model. These bounds are optimized to maintain model performance on specific datasets. To address challenge 2, upper bounds optimization can be performed using the (possibly poisoned) synthetic dataset, preserving the decentralized nature of FL. This method also minimizes impact on the model’s classification ability as it does not alter underlying parameters. Finally, to address Challenge 3, periodically adjusting these bounds is a more efficient and practical alternative to the extensive parameter tuning or trigger estimation required in other defenses.
Motivated by the above, we propose imposing an upper bound on the internal activations of the global model to constrain abnormal activations caused by backdoor threats. The defense is implemented after global model aggregation on the server and consists of three main steps: upper-bound application, optimization, and hyper-parameter adjustment, as illustrated in Fig.1 and Algorithm1. We will elaborate on the server-side defense in FL in the following sections. Note that the core FL stages, such as client updates and server-side global model fusion, remain unchanged and are not covered here. For details, please refer to [18].
@@
Step 1: Upper Bound Application. Let $\mathcal{G}(\cdot)$ denote the $L$-layer global model, derived from the model fusion at FL round $t$ [^1] For clarity, we omit $t$ in our notation.. We define the set $\mathcal{K}\subseteq\{1,\cdots,L\}$ to include all layers where bounds are applied, and the set $\mathcal{Z}=\{\mathbf{z}_{k}\in\mathbb{R}^{n_{k}}|k\in\mathcal{K}\}$ contains the upper bound vectors for activations constrain at layer $k$ within $\mathcal{K}$. Then the logit function for any class $c\in\mathcal{C}$ and any input $\mathbf{x}\in\mathbb{R}^{n_{0}}$ with activation bounds $\mathcal{Z}$ can be written as
$\overline{\mathcal{G}}_{c}(\mathbf{x;\mathcal{Z}})=\mathbf{w}_{c}^{T}(\tau_{L}\circ\sigma_{L}\circ\cdots\circ\tau_{1}\circ\sigma_{1}(\mathbf{x})),$ (1)
@@
where $\sigma_{l}:\mathbb{R}^{n_{l-1}}\to\mathbb{R}^{n_{l}}$ is the composition of the weight and activation function at the $l$-th layer of the model, $\mathbf{w}_{c}^{T}\in\mathbb{R}^{n_{L}}$ is the weight vector associated with class $c$, and $\tau_{l}(\cdot)$ applies upper bounds. For layer $l\in\mathcal{K}$, $\tau_{l}(\mathbf{v})=\min\{\mathbf{v},\mathbf{z}_{l}\}$, where the “min” operator is applied element-wise to the vectors, ensuring that the vector values are upper-bounded by $\mathbf{z}_{l}$. Otherwise, $\tau_{l}(\mathbf{v})=\mathbf{v}$, i.e., no upper bound is applied to this layer.
Step 2: Upper Bound Optimization. To ensure effective defense, activation upper bounds are minimized to allow only legitimate activations through while truncating malicious ones. A straightforward approach is to minimize the upper bounds such that the classification accuracy of the model with these bounds, evaluated on clean samples, exceeds a certain threshold. However, in the practical defense scenario considered here, the defender does not have access to clean, real data. Instead, the activation upper bounds, $\mathcal{Z}$, are optimized on the (potentially poisoned) synthetic dataset by minimizing the following Lagrangian function:
$H(\mathcal{Z},\lambda)=\sum_{\mathbf{x}\in D_{\text{syn}},c\in C}[\overline{\mathcal{G}}_{c}(\mathbf{x;\mathcal{Z}})-\mathcal{G}_{c}(\mathbf{x})]^{2}+\lambda{\sum_{l\in\mathcal{K}}{\lVert\mathbf{z}_{l}\rVert}_{2}},$ (2)
where $\overline{\mathcal{G}}_{c}(\mathbf{x;\mathcal{Z}})$ and $\mathcal{G}_{c}(\mathbf{x})$ represent the logits of the bounded and original global models, respectively.
We now elaborate on the optimization function: (i) Aligning Logits: Since the defender lacks access to clean, real data, optimizing for classification accuracy directly is not feasible. Moreover, the synthetic dataset may include mislabeled poisoned samples, making it ineffective to maximize classification accuracy on synthetic data. To mitigate the impact of mislabeled synthetic samples and ensure effective defense, the first term aligns the logits of the global model with bounds, $\overline{\mathcal{G}}_{c}(\mathbf{x;\mathcal{Z}})$, with those of the original model, $\mathcal{G}_{c}(\mathbf{x})$. This alignment ensures the model with bounds retains similar functionality on synthetic data while avoiding reliance on potentially incorrect labels. (ii) Minimizing Upper Bounds: The second term penalizes the $l_{2}-$norm of the upper bounds $\mathcal{Z}$, encouraging minimal upper bounds to effectively constrain activations. (iii) Preserving Model Parameters: To maintain the model’s original functionality and ensure the defense does not interfere with training, only the upper bounds $\mathcal{Z}$ are optimized, leaving the model parameters unchanged.
@@
Step 3: Hyper-parameter Adjustment. To ensure effective defense, the Lagrangian multiplier $\lambda$ is dynamically adjusted to balance global model performance and defense effectiveness. A hard threshold to prevent a significant drop in classification accuracy on synthetic data can undermine defense effectiveness, as discrepancies between synthetic and real data may reduce accuracy and affect the defense’s ability to constrain malicious activations. To address this, we measure the relative performance on synthetic data for hyper-parameter adjustment. Specifically, a threshold $\Delta\pi$ represents the allowable drop in classification accuracy on synthetic data. A lower $\Delta\pi$ preserves performance but may fail to adequately constrain misclassifications, whereas a higher $\Delta\pi$ enforces stricter bounds but risks overly constraining the model. The adjustment mechanism is as follows: If the accuracy drop is below $\Delta\pi$, we increase $\lambda$ by $\alpha$ to tighten the bounds; if it exceeds $\Delta\pi$, we decrease $\lambda$ to relax the constraints.
Note that after the backdoor mitigation on server, only the model parameters $\mathcal{G(\cdot)}$ are distributed to clients, excluding the upper bounds $\mathcal{Z}$. Separating the upper bounds from local training prevents compromised clients in classic attacks from adapting poisoned parameters to the bounds and thus evading the defense. The upper bounds will be distributed to the clients when the FL converges. Furthermore, fewer than five optimization iterations are sufficient to ensure defense effectiveness, with the number of optimization parameters being significantly smaller than the entire model, ensuring efficient FL training.
Input: Lagrangian multiplier $\lambda$, threshold $\Delta\pi$, factor $\alpha$
1 Initialization
2       The server queries the FM to generate $\mathcal{D}_{\text{syn}}$.
3       Randomly initialize the activation bounds $\mathcal{Z}$.
4
5for each communication round $t=1,\cdots,T$ do
6       Client Update
7             Update local model $g_{i}$ with local dataset and upload to the server.
8
9      Server Global Model Fusion
10             $\mathcal{G}\leftarrow\mathcal{E}(\{g_{i}\},\mathcal{D}_{\text{syn}})$ The aggregation function $\mathcal{E}(\cdot)$ is from [18, 13].
11
12      Server Backdoor Mitigation
13             Apply bounds $\mathcal{Z}$ to $\mathcal{G}$ to obtain $\overline{\mathcal{G}}$ (Eq.1).
14             Update $\mathcal{Z}$ by minimizing $H(\mathcal{Z},\lambda)$ (Eq.2).
15             if $\mathcal{G}(\mathcal{D}_{\text{syn}})-\overline{\mathcal{G}}(\mathcal{D}_{\text{syn}},\mathcal{Z})<\Delta\pi$ then
16                   $\lambda*=\alpha$
17
18            else
19                   $\lambda/=\alpha$
20
21             end if
22
23
24      Distribute the global model parameters $\mathcal{G}(\cdot)$ (without the activation bounds $\mathcal{Z}$) to clients.
25
26 end for
27Distribute the global model parameters with bounds $\overline{\mathcal{G}}(\cdot;\mathcal{Z})$ to clients.
Algorithm 1 Algorithm flow of the proposed defense.
IV Experiment
TABLE I: Defenses against the novel backdoor attack in homo-FL. D1: CIFAR-10, D2: CIFAR-100